OpenStudy (anonymous):

HELP! how do I solve this? the square root of x equals the square root of (x-5) +1

OpenStudy (mathmate):

Just confirming: \( \sqrt{x} = \sqrt{x-5} + 1\) Is this the equation?

OpenStudy (anonymous):

square twice and get the solution

OpenStudy (mathmate):

You can use trial and error. The answer is a small integer.

OpenStudy (anonymous):

sqrt(x)=sqrt(x-5)+1 sqrt(x)-1=sqrt(x-5) squaring both sides x+1-2*sqrt(x)=x-5 2*sqrt(x)=6 sqrt(x)=3 x=9

OpenStudy (mathmate):

Also note that the difference of two consecutive squares is an odd number: 5^2-4^2=9 4^2-3^2=7 3^2-2^2=5 2^2-1^2=3 ...

OpenStudy (anonymous):

yes mathmate thats right

OpenStudy (anonymous):

\[here x \ge 0, x \ge 5 ,so x \ge 5\] \[x=\left( x-5 \right)+2\sqrt{x-5}*1+1\] \[2\sqrt{x-5}=x-x+5-1=4\] \[\sqrt{x-5}=2\] again squaring x-5=4,x=9

OpenStudy (anonymous):

Thanks So Much!!!

OpenStudy (anonymous):

yw